# The Curse of Depth in Large Language Models

### v2

@@ A.4 Proof of Theorem 17
×$\times\Big{(}1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\ell\sigma_{x_{\ell}}}\Big{(}\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big{)}\Big{)}.$
Equation (61) is a traditional product form Whittaker and Watson (1996) for $\sigma_{x_{\ell}}$. After scaling, it becomes:
@@
The paper Vershynin (2018) tells us that:
@@
×$\quad\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
@@
×$\times\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\left(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{\frac{d_{\mathrm{head}}}{d}}\right)\right)\Big{)}\Big{)}$
×$\quad\times(\sigma_{x_{1}})^{2}\sigma\left(\sqrt{d}+\sqrt{hd_{\mathrm{head}}}\right)h$
×$\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
@@ Appendix B Variance Growth in Pre-LN Training
To examine whether the Curse of Depth also manifests in vision–language models (VLMs), we perform layer–pruning experiments on Qwen 2.5-VL-7B  (Bai et al., 2025). For both its vision encoder and language decoder, we prune one transformer layer at a time and directly evaluate the pruned model on the MMMU benchmark  (Yue et al., 2024). Figure 9 presents the resulting performance drops.
@@ Appendix D Limitations
Our study primarily focuses on Transformer-based LLMs using Pre-LN. Although Pre-LN dominates modern architectures, our theoretical study does not cover models employing alternative normalization strategies (e.g., Post-LN only (Du et al., 2021), normalization-free architectures (Zhu et al., 2025a)) or emerging paradigms such as mixture-of-experts or structured sparsity-based models.

### v3

@@ A.4 Proof of Theorem 17
$\times\Big(1+2dh\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\frac{\sigma^{2}}{\ell\sigma_{x_{\ell}}}\Big(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{d_{\mathrm{head}}/d}\right)\Big)\Big).$
Equation (61) is a traditional product form Whittaker_Watson_1996 for $\sigma_{x_{\ell}}$. After scaling, it becomes:
@@
The paper vershynin2018 tells us that:
@@
$\quad\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
@@
$\times\frac{\sigma^{2}}{\sigma_{x_{\ell}}}\left(\sigma^{2}d\sqrt{d_{\mathrm{head}}}+\left(1+\sqrt{\frac{d_{\mathrm{head}}}{d}}\right)\right)\Big)\Big)$
$\quad\times(\sigma_{x_{1}})^{2}\sigma\left(\sqrt{d}+\sqrt{hd_{\mathrm{head}}}\right)h$
$\quad\times\left[\left(\sqrt{s}+2+\frac{1}{\sqrt{s}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma_{x_{1}}\left(\sqrt{d}+\sqrt{d_{\mathrm{head}}}\right)\right].$
@@ Appendix B Variance Growth in Pre-LN Training
To examine whether the Curse of Depth also manifests in vision–language models (VLMs), we perform layer–pruning experiments on Qwen 2.5-VL-7B  (bai2025qwen2). For both its vision encoder and language decoder, we prune one transformer layer at a time and directly evaluate the pruned model on the MMMU benchmark  (yue2024mmmu). Figure 10 presents the resulting performance drops.
@@ Appendix D Limitations
Our study primarily focuses on Transformer-based LLMs using Pre-LN. Although Pre-LN dominates modern architectures, our theoretical study does not cover models employing alternative normalization strategies (e.g., Post-LN only (du2021glm), normalization-free architectures (zhu2025transformers)) or emerging paradigms such as mixture-of-experts or structured sparsity-based models.